OpenStudy (anonymous):

Solve t =p+3r/6 for r

OpenStudy (poofypenguin):

Rearrange to isolate r on one side of the equation: t =p+3r/6 t-p=3r/6 6*(t-p)=3r [6*(t-p)]/3=r

OpenStudy (anonymous):

Whats next?

OpenStudy (poofypenguin):

that's it, you can't really do any more without more given information.